(include formulas)

1. Suppose you want to have $400,000 for retirement in 20 years. Your account earns 4% interest. How much would you need to deposit in the account each month?

2. You have $500,000 saved for retirement. Your account earns 10% interest. How much will you be able to pull out each month, if you want to be able to take withdrawals for 25 years?
Answer #1

Ans 1) $ 1090.59

P = Periodic payments
r = rate of interest
n = no of years
Future Value of Annuity = P ( (1 + r)n - 1 ) / r
400000 = P* ((1 + 4%/12)^240 - 1) / (4%/12)
400000 = 366.774626089931 * P
P = 400000 / 366.774626089931
P = 1090.59